OpenStudy (anonymous):

solve dy/dx=1/(1+2x*cot(y))

OpenStudy (shayaan_mustafa):

integrate with respect to x. \[\LARGE \int\limits\frac{dy}{dx}dx=\int\limits\frac{1}{1+2x*\cot(y)}dx\]

OpenStudy (anonymous):

is it y=0??